# Robust Learning of Diverse Code Edits

### v1

@@ Abstract
Software engineering activities frequently involve edits to existing code. However, contemporary code language models (LMs) lack the ability to handle diverse types of code-edit requirements. In this work, we attempt to overcome this shortcoming through (1) a novel synthetic data generation pipeline and (2) a robust model adaptation algorithm. Starting with seed code examples and diverse editing criteria, our pipeline generates high-quality samples comprising original and modified code, along with natural language instructions in different styles and verbosity. Today’s code LMs come bundled with strong abilities, such as code generation and instruction following, which should not be lost due to fine-tuning. To ensure this, we propose a novel adaptation algorithm, SeleKT, that (a) leverages a dense gradient-based step to identify the weights that are most important for code editing, and (b) does a sparse projection onto the base model to avoid overfitting. Using our approach, we obtain a new series of models NextCoder (adapted from QwenCoder-2.5) that achieves strong results on five code-editing benchmarks, outperforming comparable size models and even several larger ones. We show the generality of our approach on two model families (DeepSeekCoder and QwenCoder), compare against other fine-tuning approaches, and demonstrate robustness by showing retention of code generation abilities post adaptation.
11footnotetext: *equal contribution.22footnotetext: Corresponding authors: tushar.aggarwal53@gmail.com, {t-swsingh,abawasthi,kanadeaditya,nagarajn}@microsoft.com
@@ 3 Synthesizing a Diverse Code Editing Dataset
Thus, in addition to using CommitPackFT as a source of supervision, we propose a novel method for generating high-quality code-editing examples starting with real seed data from GitHub using large and medium LMs, such as GPT-4o and the instruct version of Llama-3.3-70B. Our method provides greater diversity by offering explicit control over varying levels of granularity (function, class, and file-level code edits), a wide range of code edit types (e.g., bug fixing, latency and runtime improvements, addressing security vulnerabilities, optimizing resource utilization, and enhancing maintainability), diversity in programming languages, and varying levels of instruction complexity (both concise and verbose, and single-turn as well as multi-turn style conversations). Table 1 provides statistics of the code editing data generated using our method across eight programming languages. We generate approximately 127K examples and 229M tokens, with 48K examples produced using GPT-4o.
@@
Finally, to ensure high quality of generated examples, we prompt the LLM to verify whether the target file is a correctly edited version of the source file, consistent with the instructions generated in the previous step. The model evaluates the instance by assigning a score from 0 to 10 across five criteria: (1) Correctness of edits w.r.t. requested improvements (e.g. latency), (2) adherence of the edits to the instructions, (3) code quality, (4) instruction quality, and (5) the usefulness of the example for fine-tuning small models. An instance is deemed valid if its average score is at least 7, and all individual scores exceed 5. The prompt used for this step is given in Figure 8.
@@ 4 Robust Model Adaptation
While the projection step itself is straight-forward, computation of dense gradients can be more expensive. The cost is mitigated to some extent by doing mini-batching and restricting the total number of epochs to a few. We show in Sections 5.2 and 5.3 that the resulting improvements in generalization accuracy are significant.
@@ 5.1 Experimental Setup
Hyperparameters  (i) We used a temperature of 0.6 for data generation for both Llama-3.3-70B and GPT-4o. (ii) We fix the periodicity to 1 epoch in the SeleKT algorithm unless specified otherwise, i.e., $M=$ total number of mini-batches. We set sparsity $\alpha=0.05$ per layer. We selected these values based on initial experiments. In Section 5.7, we show ablations on these choices.
@@
Table 3: Evaluation datasets: We use five diverse code editing benchmarks with varying input granularity and code-editing criteria. Additionally, we also evaluate on two code generation tasks.
@@
Additionally, we also evaluate the models on standard code generation benchmarks, HumanEval+ and MBPP+ [24] (Table 3, rows 6–7), to measure the extent to which models fine-tuned for code editing still retain their code generation and instruction following abilities. These benchmarks require a model to complete a function given its signature and a short description. The evaluation metric is execution accuracy on test cases. Appendix A.1.3 provides additional details on the benchmarks.
@@ 5.4 Effectiveness of Synthetic Data
Table 6 compares the performance of DeepSeekCoder-6.7B fine-tuned separately on CommitPackFT and on the synthetic data generated by our pipeline. While synthetic data offers a performance advantage over CommitPackFT, the latter represents a high-quality sample of real-world developer commits. To enhance generalizability, we incorporate both datasets in our fine-tuning process.
@@
5.5 Effectiveness of SeleKT on Different Model Sizes
Models HumanEvalFix CanItEdit Aider
QwenCoder-2.5-3B 73.2 37.1 36.8
NextCoder-3B 75.6 42.4 37.6
QwenCoder-2.5-14B 87.8 58.1 66.9
NextCoder-14B 89.8 60.2 72.2
QwenCoder-2.5-32B 90.2 61.0 72.9
NextCoder-32B 88.9 62.4 74.7
Table 7 demonstrates the performance of our SeleKT algorithm across various model sizes. For the smaller 3B model, NextCoder-3B consistently improves performance across all three benchmarks, with particularly notable gains on the CanItEdit benchmark (+5.3%). The mid-sized 14B model shows similar improvements, with SeleKT enhancing performance across all datasets, achieving the most substantial improvement on the Aider benchmark (+5.3%). For the largest 32B model, there is a slight decrease in HumanEvalFix performance (-1.3%), but noticeable gains in the CanItEdit and Aider benchmarks. These results suggest that our training pipeline provides consistent gains across model sizes.
@@ 6 Conclusions
The ability of code LMs to accurately edit code at different scales and based on diverse instructions is central to their use in software engineering. In this paper, we present the next step in enhancing the code-editing ability of code LMs by developing a synthetic data pipeline and a robust adaptation algorithm SeleKT. Our pipeline produces diverse data that improves model performance and the adaptation algorithm ensures that the general, pre-learned abilities of the models are not lost during fine-tuning. We comprehensively evaluate our method on multiple code editing and generation benchmarks to establish these claims. In future, we want to extend our pipeline to cover more scenarios and evaluate SeleKT in tasks other than code-editing.
@@ A.1.3 Benchmarks
A.2 Prompts
@@ Appendix C Examples Showing the Quality of Our Model Outputs
Figure 10: Example code-editing task (top) from Aider benchmark. NextCoder-7B (RHS) solves the task correctly in the first attempt, while QwenCoder-2.5-7B (LHS) gets it wrong in both the attempts: The directions are represented by pairs of the form (a,b) where a is the y-coordinate and b is the x-coordinate. However, QwenCoder-2.5-7B incorrectly assumes that a and b are respectively x and y coordinates and generates edited code accordingly. NextCoder-7B interprets the instruction accurately and generates correct code. The relevant code in NextCoder-7B’s output is highlighted in bold.
Figure 11: Example code-editing task (top) from Aider benchmark. NextCoder-7B (RHS) solves the task correctly in the first attempt: The input can only contain spaces or ‘*’s. However, in the example in the prompt, ’.’ is used to represent space and the same is explicitly stated. Despite this, QwenCoder-2.5-7B accepts inputs which contain characters other than ’*’ and space. This problem is fixed by NextCoder-7B by rejecting any input which violates this constraint. The relevant code in NextCoder-7B’s output is highlighted in bold.

### v2

@@ Abstract
Software engineering activities frequently involve edits to existing code. However, contemporary code language models (LMs) lack the ability to handle diverse types of code-edit requirements. In this work, we attempt to overcome this shortcoming through (1) a novel synthetic data generation pipeline and (2) a robust model adaptation algorithm. Starting with seed code examples and diverse editing criteria, our pipeline generates high-quality samples comprising original and modified code, along with natural language instructions in different styles and verbosity. Today’s code LMs come bundled with strong abilities, such as code generation and instruction following, which should not be lost due to fine-tuning. To ensure this, we propose a novel adaptation algorithm, SeleKT, that (a) leverages a dense gradient-based step to identify the weights that are most important for code editing, and (b) does a sparse projection onto the base model to avoid overfitting. Using our approach, we obtain a new series of models NextCoder (adapted from QwenCoder-2.5) that achieves strong results on five code-editing benchmarks, outperforming comparable size models and even several larger ones. We show the generality of our approach on two model families (DeepSeekCoder and QwenCoder), compare against other fine-tuning approaches, and demonstrate robustness by showing retention of code generation and general problem-solving abilities post adaptation. We opensource the models, synthetic dataset, and implementation at aka.ms/nextcoder.
11footnotetext: *equal contribution.22footnotetext: Corresponding authors: tushar.aggarwal53@gmail.com, {t-swsingh,abawasthi,kanadeaditya,nagarajn}@microsoft.com33footnotetext: To appear in ICML 2025 as “NextCoder: Robust Adaptation of Code LMs to Diverse Code Edits”.
@@ 1 Introduction
4. Opensource: We opensource the models, synthetic dataset, and implementation at aka.ms/nextcoder.
@@ 3 Synthesizing a Diverse Code Editing Dataset
Thus, in addition to using CommitPackFT as a source of supervision, we propose a novel method for generating high-quality code-editing examples starting with real seed data from GitHub using large and medium LMs, such as GPT-4o and the instruct version of Llama-3.3-70B. Our method provides greater diversity by offering explicit control over varying levels of granularity (function, class, and file-level code edits), a wide range of code edit types (e.g., bug fixing, latency and runtime improvements, addressing security vulnerabilities, optimizing resource utilization, and enhancing maintainability), diversity in programming languages, and varying levels of instruction complexity (both concise and verbose, and single-turn as well as multi-turn style conversations). Table 1 provides statistics of the code editing data generated using our method across eight programming languages. We generate approximately 127K examples and 229M tokens, with 48K examples produced using GPT-4o. We perform contamination analysis of our generated data with respect to the evaluation benchmarks in Section A.1.5.
@@
Finally, to ensure high quality of generated examples, we prompt the LLM to verify whether the target file is a correctly edited version of the source file, consistent with the instructions generated in the previous step. The model evaluates the instance by assigning a score from 0 to 10 across five criteria: (1) Correctness of edits w.r.t. requested improvements (e.g. latency), (2) adherence of the edits to the instructions, (3) code quality, (4) instruction quality, and (5) the usefulness of the example for fine-tuning small models. An instance is deemed valid if its average score is at least 7, and all individual scores exceed 5. The prompt used for this step is given in Figure 8. We also performed a human study to assess the quality of the generated data, and the details are in Section A.3.
@@ 4 Robust Model Adaptation
While the projection step itself is straight-forward, computation of dense gradients can be more expensive. The cost is mitigated to some extent by doing mini-batching and restricting the total number of epochs to a few. We show in Sections 5.2 and 5.4 that the resulting improvements in generalization accuracy are significant.
@@ 5.1 Experimental Setup
Hyperparameters  (i) We used a temperature of 0.6 for data generation for both Llama-3.3-70B and GPT-4o. (ii) We fix the periodicity to 1 epoch in the SeleKT algorithm unless specified otherwise, i.e., $M=$ total number of mini-batches. We set sparsity $\alpha=0.05$ per layer. We selected these values based on initial experiments. In Section 5.7, we show ablations on these choices. Additional hyperparameter tuning for SFT and LoRA are detailed in Section A.1.4.
@@
GSM8K Mathematical problem solving 1.32K
MMLU Multiple-choice STEM knowledge 3.15K
Table 3: Evaluation datasets: We use four diverse code editing benchmarks with varying input granularity and code-editing criteria. Additionally, we also evaluate on two code generation tasks and two non-coding tasks.
@@
Additionally, we also evaluate the models on (a) standard code generation benchmarks, HumanEval+ and MBPP+ [26] (Table 3, rows 6–7), to measure the extent to which models fine-tuned for code editing still retain their code generation abilities. These benchmarks require a model to complete a function given its signature and a short description. The evaluation metric is execution accuracy on test cases. Further, we do evaluation on (b) non-coding benchmarks, GSM8K [4] and STEM subset of MMLU [12] (Table 3, rows 8–9), to gauge how well they maintain math problem solving and natural-language understanding capabilities. Appendix A.1.3 provides details on the benchmarks.
@@ 5.2 Performance on Code-Editing Benchmarks
5.3 Effectiveness of SeleKT on Different Model Sizes
Models HumanEvalFix CanItEdit Aider Aider Polyglot
QwenCoder-2.5-3B 73.2 37.1 36.8 -
QwenCoder-2.5-3B-LoRA 64.6 36.2 35.8 -
QwenCoder-2.5-3B-SFT 76.2 32.4 30.1 -
NextCoder-3B 75.6 42.4 37.6 -
QwenCoder-2.5-14B 87.8 58.1 66.9 9.3
QwenCoder-2.5-14B-LoRA 78.0 50.9 66.2 5.3
QwenCoder-2.5-14B-SFT 79.9 42.4 36.8 3.1
NextCoder-14B 89.8 60.2 72.2 12.2
QwenCoder-2.5-32B 90.2 61.0 72.9 16.4
QwenCoder-2.5-32B-LoRA 82.3 52.4 60.2 6.7
QwenCoder-2.5-32B-SFT 81.7 49.5 66.9 8.4
NextCoder-32B 88.9 62.4 74.7 21.9
Table 5 demonstrates the performance of our SeleKT algorithm across various model sizes, including multilingual capabilities measured by the Aider Polyglot benchmark. For the smaller 3B model, NextCoder-3B shows significant improvements over the base model across most benchmarks, with a substantial gain on the CanItEdit benchmark (+5.3%). Notably, while the QwenCoder-2.5-3B-SFT achieves slightly better performance on HumanEvalFix, our approach excels on other benchmarks. At the 14B scale, NextCoder-14B consistently outperforms all baseline variants, achieving gains across all four benchmarks, with particularly impressive improvements on the Aider Polyglot benchmark (+2.9% over base model). For the largest 32B model, while there is a slight decrease in HumanEvalFix performance compared to the base model (-1.3%), NextCoder-32B achieves the highest scores across all other benchmarks, with a remarkable improvement on Aider Polyglot (+5.5%). These results demonstrate that our training approach provides consistent gains across model sizes.
@@ 5.4 Preserving Pre-learned Knowledge
Model MMLU GSM8K
QwenCoder-2.5-7B 53.0 83.40
QwenCoder-2.5-32B 71.9 93.71
NextCoder-7B 54.5 81.65
NextCoder-32B 72.7 92.65
Table 7: Accuracy (%) on MMLU and GSM8K benchmarks.
The results in Table 7 demonstrate that NextCoder-32B improves on MMLU scores (+0.8 points), but has a small regression on GSM8K (-1.06 points). Overall, these results demonstrate that our training methodology induces specialized code editing capabilities without sacrificing math problem solving and natural-language understanding abilities.
@@ 5.5 Effectiveness of Synthetic Data
Table 8 compares the performance of DeepSeekCoder-6.7B fine-tuned separately on CommitPackFT and on the synthetic data generated by our pipeline. While synthetic data offers a performance advantage over CommitPackFT, the latter represents a high-quality sample of real-world developer commits. To enhance generalizability, we incorporate both datasets in our fine-tuning process. Performance variations across different data sizes are further analyzed in Section A.2.
@@ 6 Conclusions
The ability of code LMs to accurately edit code at different scales and based on diverse instructions is central to their use in software engineering. In this paper, we present the next step in enhancing the code-editing ability of code LMs by developing a synthetic data pipeline and a robust adaptation algorithm SeleKT. Our pipeline produces diverse data that improves model performance and the adaptation algorithm ensures that the general, pre-learned abilities of the models are not lost during fine-tuning. We comprehensively evaluate our method on multiple code editing and generation benchmarks to establish these claims. In future, we want to extend our pipeline to cover more scenarios and evaluate SeleKT in tasks other than code-editing such as mathematical reasoning and natural language tasks.
@@ A.1.3 Benchmarks
Additionally, we evaluate the models on non-coding benchmarks.
MMLU benchmark [12] (STEM subset) comprises 3.15K problems spanning Physics, Chemistry, Biology, Computer Science, Mathematics, and Engineering. We employed the few-shot setting (N=4) with the prompt configuration from Qwen models’ official evaluation script.
GSM8K benchmark [4] contains grade-school mathematical problems that test reasoning abilities. For evaluation, we used the few-shot setting (N=4) with the prompt configuration from Qwen models’ official evaluation script.
A.1.4 Hyperparameter tuning for Baselines
Method CanItEdit HumanEvalFix Aider
SFT 46.2 78.6 52.6
LoRA 43.8 79.9 54.1
SeleKT 50.48 81.10 65.70
Table A.2: Results for the best versions of LoRA and SFT for QwenCoder-2.5-7B
To ensure fair comparison between our proposed method and baseline approaches, we conducted systematic hyperparameter tuning experiments for SFT and LoRA methods. Since SeleKT was optimized with a learning rate of 1e-5 and weight decay of 0.0, we explored comparable configurations for the baseline methods. For both SFT and LoRA, we investigated learning rates of 2e-6 and 5e-6 combined with weight decay values of 0.10 and 0.05. For LoRA specifically, we further evaluated rank values ranging from 16 to 64 and alpha parameters of 8 and 16.
While this comprehensive tuning process yielded performance improvements for both baseline methods, Table A.2 demonstrates that a substantial performance gap persists between these optimized baselines and our NextCoder-7B model. Across all three evaluation benchmarks (CanIEdit, HumanEvalEdit, and Aider), SeleKT consistently outperforms the tuned baseline approaches. The most pronounced difference appears in the Aider benchmark, where SeleKT achieves a 11.6% improvement over the best baseline. These results underscore the effectiveness of our selective parameter updating approach beyond what can be achieved through standard hyperparameter optimization of conventional fine-tuning methods.
A.1.5 Decontaminating Synthetic Training Data
To ensure the integrity of our evaluations, we adopted the decontamination procedure used in prior works such as StarCoder [25, 28]. Specifically, we applied near-duplicate detection using MinHash and Locality Sensitive Hashing (LSH) to identify and remove any potential overlaps between our training dataset and evaluation benchmarks (CanItEdit, HumanEvalFix, Aider, NoFunEval). Our decontamination analysis confirmed 0% data leakage, verifying that no benchmark data was present in the training set.
A.2 Robustness of SeleKT across dataset sizes
Dataset size CanItEdit HumanEvalFix Aider
Base model (QwenCoder-2.5-7B) 48.1 73.8 59.4
25% 47.67 80.20 60.90
50% 48.57 81.43 62.70
75% 49.01 81.02 63.80
100% (NextCoder-7B) 50.48 81.10 65.70
Table A.3: Performance of QwenCoder-2.5-7B when finetuned on splits of different sizes of our dataset
To assess scalability w.r.t. to training data size, we finetuned the QwenCoder-2.5-7B model on varying fractions (random sampling 25%, 50% and 75%) of our dataset which includes both synthetic and CommitPackFT data. The results are presented in the Table A.3. All models were trained for 3 epochs. The results show a clear trend: while performance on CanItEdit and Aider sees a drop at 25% w.r.t to the base model, increasing the dataset size consistently improves performance across all benchmarks (CanItEdit, HumanEvalFix, and Aider).
A.3 Human Evaluation of Synthetic Dataset
We conducted a pilot human-study to assess the quality of the generated training dataset. In this study, we involved three participants who have 3-4 years of experience in software development with strong expertise in Python. We randomly selected 100 samples from the Python split of our synthetic dataset and asked participants to rate each sample on the scale of 1-5 (1 being poor quality and 5 being excellent quality) on the following three questions:
i) Instruction Usefulness:
How well does the detailed instruction capture a potential code-editing scenario with respect to the original code?
ii) Instruction Consistency:
How consistent are the three styles of instructions (detailed, concise and human-like) with each other and with the respective styles?
iii) Solution Correctness:
How well does the edited code match the edit described in the detailed instruction?
Metric Participant 1 Participant 2 Participant 3 Overall Mean Overall SD
Instruction Usefulness 4.92 ± 0.27 4.93 ± 0.26 4.38 ± 0.60 4.74 0.48
Instruction Consistency 4.29 ± 0.48 4.88 ± 0.32 4.55 ± 0.54 4.57 0.51
Solution Correctness 4.92 ± 0.27 4.96 ± 0.24 4.60 ± 0.51 4.83 0.40
Overall Mean ± SD 4.71 ± 0.46 4.92 ± 0.28 4.51 ± 0.56
Table A.4: Evaluation of Instruction Usefulness, Consistency, and Solution Correctness
In the table A.4, we present the mean (along with standard deviations) ratings by participant and by question. The scores are consistently close to the highest score of 5 across all participants and questions. This provides a strong indication of human-machine alignment, with low to moderate variance. This study helps validate that our synthetic data generation pipeline is able to generate samples that meet human expectations in terms of quality and consistency.
Score (Higher is better) Instruction Usefulness Instruction Consistency Solution Correctness
5 229 175 250
4 65 122 48
3 6 3 2
Total 300 300 300
Table A.5: Distribution of Evaluation Scores
In the table A.5, we have shown the scores distribution. Clearly, all samples received scores 3 (neutral quality) or above on all the questions. We particularly inspected the samples that received the neutral rating (score 3) since those were perceived as relatively low-quality samples by one or more participants. We made the following observations:
i) Instruction-Edit Misalignment:
In some cases, instructions correctly described the intent but the edits were not entirely appropriate. For example, in response to an instruction to handle datetime parsing, the edited code parsed dates against raw strings, which would cause runtime errors.
ii) Incomplete Error Handling:
Some examples did introduce error handling, but overlooked edge cases (e.g., what if the ‘tasks.json‘ file exists but is empty?).
iii) Style Inconsistency:
A few participants noted that stylistic or structural variations across instruction formats led to minor misunderstandings of the code-editing intent.
A.4 Prompts
@@ Appendix C Examples Showing the Quality of Our Model Outputs
Figure 10: Example code-editing task (top) from Aider benchmark. NextCoder (RHS) solves the task correctly in the first attempt, while QwenCoder-2.5-7B (LHS) gets it wrong in both the attempts: The directions are represented by pairs of the form (a,b) where a is the y-coordinate and b is the x-coordinate. However, QwenCoder-2.5-7B incorrectly assumes that a and b are respectively x and y coordinates and generates edited code accordingly. NextCoder interprets the instruction accurately and generates correct code. The relevant code in NextCoder’s output is highlighted in bold.
Figure 11: Example code-editing task (top) from Aider benchmark. NextCoder (RHS) solves the task correctly in the first attempt: The input can only contain spaces or ‘*’s. However, in the example in the prompt, ’.’ is used to represent space and the same is explicitly stated. Despite this, QwenCoder-2.5-7B accepts inputs which contain characters other than ’*’ and space. This problem is fixed by NextCoder by rejecting any input which violates this constraint. The relevant code in NextCoder’s output is highlighted in bold.